# Tip of the Tongue Query Elicitation for Simulated Evaluation

### v1

@@ Abstract.
Tip-of-the-Tongue Known-Item Retrieval, Synthetic Query Generation, Human Query Elicitation
Equal Contribution.
Now at Meta.
copyright: rightsretainedcopyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/2025/06ccs: Information systems Test collectionsccs: Information systems Specialized information retrievalccs: Human-centered computing Interaction design
@@ 1. Introduction
The evaluation of TOT retrieval systems faces significant challenges due to limitations in test collection resources and current dataset construction methods. Most existing TOT queries in current datasets are sourced from community question answering (CQA) websites, which fail to fully capture the diversity of real-world TOT information needs (Meier et al., 2021). This reliance on CQA platforms introduces domain skewness, as existing datasets are heavily biased toward casual leisure topics like movies and books (Elsweiler et al., 2011b), while underrepresenting areas such as people and landmarks. Even attempts to expand domain coverage (Meier et al., 2021) have had limited success, leaving many domains underrepresented. Additionally, the labor-intensive nature of dataset construction—which relies on manual annotation, coding, and refinement (Bhargav et al., 2022, 2023; Arguello et al., 2021; Meier et al., 2021)—slows progress and restricts the ability to conduct large-scale and multi-domain evaluations.
@@ 2. Related Work
2.1. Query Simulation and Know-Item Retrieval
Query simulation methods have been used for various purposes, including document expansion (Nogueira et al., 2019) and synthetic test collection generation (Rahmani et al., 2024). In the context of known-item retrieval, these methods have been explored to improve retrieval strategies (Ogilvie and Callan, 2003) and evaluation frameworks (Azzopardi and de Rijke, 2006; Hagen et al., 2015).
@@ 3.2.1. System Rank Correlation
Figure 1. Validation of elicited queries using system rank correlation. We evaluate 40 different retrieval models using both CQA-based and elicited queries, ranking them based on search performance measured by MRR and NDCG. We then compute Kendall’s Tau and Pearson correlation to assess the agreement between the rankings derived from the two query sets.
@@ 3.2.2. Linguistic Similarity
To conduct this analysis, we annotate our elicited TOT queries in the Movie domain using these linguistic codes and compare the percentage distributions of codes found in CQA-based and elicited queries. To automate this process, we develop a language model-based automatic code annotator, prompting GPT-4o-mini [^3] Temperature is set to 0 for reproducibility and consistency. to produce JSON-formatted sentence-level annotations.
@@ 4.1. Synthetic Query Generation
    The entity’s Wikipedia page is retrieved and used to construct a summarization prompt for the GPT-4o model [^4] GPT4o-2024-05-13. To ensure the content fits within the model’s token limit, the page is truncated if necessary. The model processes the page into a two-paragraph summary. For Movie domain entities, we check for a dedicated Plot section and ensure its content is included in the summarization prompt, as it often provides essential context for generating realistic TOT queries.
@@ 4.2. Results and Analyses
We tested various configurations to identify the most effective TOT query generation approach. The final prompt structure was determined through the extensive testing of different prompt templates, system roles, instruction types, and temperature settings, as well as evaluating the inclusion of Wikipedia summaries (Table 1). For each configuration, queries were generated using entities from the test set of the TREC TOT Track 2023666https://trec-tot.github.io/guidelines-2023, derived from the MS-TOT dataset (Arguello et al., 2021), and evaluated based on system rank correlation and linguistic similarity.
@@
Based on findings across three validation methods, we empirically conclude that LLM-generated TOT queries are an effective approach for designing simulated evaluations of TOT retrieval systems.
@@
To validate the generated queries, we compared them against CQA-collected queries. Since existing TOT test collections are largely focused on the Movie domain, no established test collection exists for the Landmark and Person domains. To address this, we hand-curated a test collection, resulting in 22 Landmark-domain queries and 70 Person-domain queries from the Reddit /r/tipofmytongue subreddit.
@@ 5.1. Visual Stimuli Collection
To obtain images in Movie, Landmark, and Person domains, we collected entities from Wikipedia, selecting pages based on their infobox namespace [^7] https://en.wikipedia.org/wiki/Wikipedia:List_of_infoboxes. and the presence of images. For the Person and Landmark domains, we directly collected images from Wikipedia pages. However, for the Movie domain, we identified linked IMDb or TMDB IDs on the Wikipedia pages and retrieved high-quality backdrop or still-cut images from TMDB. This approach was necessary because Wikipedia’s movie images were predominantly posters, which often contain the movie title and cannot be used to elicit a TOT state. TMDB backdrops, on the other hand, provide spoiler-free, high-quality visuals suitable for this purpose. [^8] https://www.themoviedb.org/bible/image/ For the Landmark domain, we selected “Place” as a top-level category, and for the Person domain, we chose “Person”.
@@
Next, we collected Wikipedia pages associated with the selected infobox namespaces and sorted them by page view counts, using the Wikipedia database dump from October 1, 2023.999https://dumps.wikimedia.org/enwiki/ We selected the top 20% of pages by view count, as less popular entities were deemed unlikely to be recognized by participants.
@@ 5.2. Interface Design
Figure 5. Flowchart of the Human TOT Query Elicitation Interface.
@@ 6. Discussion
The full release of these queries is scheduled for the TREC 2025 TOT track, where they will be included as part of the official test collection. However, for TREC 2024, and in this work, we have released 450 synthetic queries (150 per domain) from the full set of generated queries. Each query in the dataset is accompanied by its corresponding Wikidata ID, domain name, and entity name, ensuring clear entity association for retrieval experiments and analysis. Alongside these queries, we provide the source code for query generation and experimentation, as well as the visual stimuli entity set with corresponding image URLs and Wikidata ID. We also release the MTurk-based human query collection interface, allowing researchers to replicate or extend the human TOT query elicitation process.
Availability. At the time of review, LLM-elicited queries are publicly available as part of the TREC 2024 TOT track test collection [^12] https://github.com/kimdanny/llm-tot-query-elicitation and can also be accessed at the track website [^13] https://trec-tot.github.io/guidelines-2024. The human-elicited queries will be released as part of the TREC 2025 TOT track, aligning with the SIGIR 2025 conference. Although the human-elicited queries are not yet publicly available, we have released the source code for the human query collection interface used in pilot testing on Amazon MTurk [^14] https://github.com/kimdanny/human-tot-query-elicitation-mturk. This allows researchers to explore and reproduce the query elicitation process for future development.
@@
Limitation and Future Work. A limitation of our current LLM-elicitation method is that prompts are domain-specific, limiting their generalizability. Future work should develop generalized prompting strategies to elicit TOT queries across diverse search contexts without extensive manual tuning. Additionally, expanding the methodology to multi-turn interactions could better reflect real-world TOT search behavior, where users iteratively refine queries as they recover missing information. Simulating this step-by-step recall process could improve query realism and retrieval effectiveness, further advancing TOT retrieval research.
@@ 7. Conclusion
8. Acknowledgment
We thank Ian Soboroff for implementing the in-house human query collection interface based on the MTurk pilot version, as well as for coordinating the hiring of human contractors and overseeing the collection of human TOT queries at NIST. We also appreciate Alfredo Gomez for assisting with the manual collection of validation queries for the Landmark and Person domains. Additionally, we thank Alfredo Gomez, Athiya Deviyani, Jessica Huynh, and Shaily Bhatt for their valuable feedback during the pilot testing of the human TOT query collection interface.
@@ Appendix C Prompt for Linguistic Similarity Test in the Movie Domain
Given a paragraph, you will annotate the paragraph with the following coding scheme. Read the paragraph and annotate each sentence following the coding scheme. When outputting the response, follow the output JSON format. Coding scheme: movie: Sentence that describes characteristics of the movie itself (e.g., category, genre, tone, plot). context: Context where the searcher previously engaged with the movie. References to when, where the searcher watched the movie, who they watched it with, and even world events that were happening around the time. previous-search: Describes a previous attempt to find the movie title. References to failed search attempts, descriptions of search strategies that were unsuccessful. social: Communicates a social nicety to the community of people (e.g., If you could help I’d really appreciate it). uncertainty: Conveys uncertainty about information described (e.g., It was a foreign film I think either French or German, but I could be wrong). opinion: Conveys an opinion or judgment about some aspect of the movie (e.g., Its pretty confusing all the way to the end when there’s only one surviving woman). emotion: Conveys how the movie made the viewer feel (e.g., It was the first movie that kept me awake at night). relative-comparison: Statement that draws one or more comparisons that require additional information in order to extract the meaning of the statement (e.g., the main character looks like Brad Pitt). Paragraph: {Paragraph} Output JSON format: { ”¡Sentence Number¿”: [”¡code¿”], ”¡Sentence Number¿”: [”¡code¿”, ”¡code¿”, …], ”¡Sentence Number¿”: [], } For each sentence number in a paragraph, you can put multiple appropirate code. However, if none of the codes match with a sentence, you can leave the list empty.

### v2

@@ Abstract.
Tip-of-the-Tongue Known-Item Retrieval; Synthetic Query Generation; Human Query Elicitation
copyright: rightsretainedjournalyear: 2025copyright: ccconference: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 13–18, 2025; Padua, Italybooktitle: Proceedings of the 48th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’25), July 13–18, 2025, Padua, Italydoi: 10.1145/3726302.3730335isbn: 979-8-4007-1592-1/2025/07ccs: Information systems Test collectionsccs: Information systems Specialized information retrievalccs: Human-centered computing Interaction design
@@ 1. Introduction
The evaluation of TOT retrieval systems faces significant challenges due to limitations in test collection resources and current dataset construction methods. Most existing TOT queries in current datasets are sourced from community question answering (CQA) websites, which fail to fully capture the diversity of real-world TOT information needs (Meier et al., 2021). This reliance on CQA platforms introduces domain skewness, as existing datasets and studies are heavily biased toward casual leisure topics like movies and books (Elsweiler et al., 2011b; Bogers et al., 2025), while underrepresenting areas such as people and landmarks. Even attempts to expand domain coverage (Meier et al., 2021) have had limited success, leaving many domains underrepresented. Additionally, the labor-intensive nature of dataset construction—which relies on manual annotation, coding, and refinement (Bhargav et al., 2022, 2023; Arguello et al., 2021; Meier et al., 2021)—slows progress and restricts the ability to conduct large-scale and multi-domain evaluations.
@@ 2. Related Work
2.1. Query Simulation & Known-Item Retrieval
Query simulation methods have been used for various purposes, including document expansion (Nogueira et al., 2019) and synthetic test collection generation (Rahmani et al., 2024). In the context of known-item retrieval, these methods have been explored to improve retrieval strategies (Ogilvie and Callan, 2003; Phanse et al., 2024) and evaluation frameworks (Azzopardi and de Rijke, 2006; Hagen et al., 2015).
@@ 3.2.1. System Rank Correlation
Figure 1. Validation of elicited queries using system rank correlation. We evaluate 40 different retrieval models using both CQA-based and elicited queries, ranking them based on search performance measured by MRR@1000 and NDCG@1000. We then compute Kendall’s Tau and Pearson correlation to assess the agreement between the rankings derived from the two query sets.
@@ 3.2.2. Linguistic Similarity
To conduct this analysis, we annotate our elicited TOT queries in the Movie domain using these linguistic codes and compare the percentage distributions of codes found in CQA-based and elicited queries. To automate this process, we develop a language model-based automatic code annotator, prompting GPT-4o-mini to produce JSON-formatted sentence-level annotations. [^3] GPT model’s temperature is set to 0 for reproducibility and consistency.
@@ 4.1. Synthetic Query Generation
    The entity’s Wikipedia page is retrieved and used to construct a summarization prompt for the GPT-4o model. [^4] GPT4o-2024-05-13 To ensure the content fits within the model’s token limit, the page is truncated if necessary. The model processes the page into a two-paragraph summary. For Movie domain entities, we check for a dedicated Plot section and ensure its content is included in the summarization prompt, as it often provides essential context for generating realistic TOT queries.
@@ 4.2. Results and Analyses
We tested various configurations to identify the most effective TOT query generation approach. The final prompt structure was determined through the extensive testing of different prompt templates, system roles, instruction types, and temperature settings, as well as evaluating the inclusion of Wikipedia summaries (Table 1). For each configuration, queries were generated using entities from the test set of the TREC TOT Track 2023 (Arguello et al., 2023), derived from the MS-TOT dataset (Arguello et al., 2021), and evaluated based on system rank correlation and linguistic similarity.
@@
Additionally, linguistic similarity analysis using Earth Mover’s Distance (EMD) resulted in a value of 0.0264 (Figure 4) in Experiment 13, showing the lowest value.
Based on findings across the two validation methods, we empirically conclude that LLM-generated TOT queries are an effective approach for designing simulated evaluations of TOT retrieval systems.
@@
To validate the generated queries, we compared them against CQA-collected queries. Since existing TOT test collections are largely focused on the Movie domain, no established test collection exists for the Landmark and Person domains. To address this, we hand-curated a test collection, resulting in 22 Landmark-domain queries and 70 Person-domain queries from the Reddit’s subreddit, called /r/tipofmytongue.
@@ 5.1. Visual Stimuli Collection
To obtain images in Movie, Landmark, and Person domains, we collected entities from Wikipedia, selecting pages based on their infobox namespace and the presence of images. [^6] List of infoboxes: https://en.wikipedia.org/wiki/Wikipedia:List_of_infoboxes. For the Person and Landmark domains, we directly collected images from Wikipedia pages. However, for the Movie domain, we identified linked IMDb or TMDB IDs on the Wikipedia pages and retrieved high-quality backdrop or still-cut images from TMDB. This approach was necessary because Wikipedia’s movie images were predominantly posters, which often contain the movie title and cannot be used to elicit a TOT state. TMDB backdrops, on the other hand, provide spoiler-free, high-quality visuals suitable for this purpose. [^7] https://www.themoviedb.org/bible/image/ For the Landmark domain, we selected “Place” as a top-level category, and for the Person domain, we chose “Person”.
@@
Next, we collected Wikipedia pages associated with the selected infobox namespaces and sorted them by page view counts, using the Wikipedia database dump from October 1, 2023.888https://dumps.wikimedia.org/enwiki/ We selected the top 20% of pages by view count, as less popular entities were deemed unlikely to be recognized by participants.
@@
Figure 5. Flowchart of the Human TOT Query Elicitation Interface.
@@ 6. Discussion
The full release of these queries is scheduled for the TREC 2025 TOT track, where they will be included as part of the official test collection. However, for TREC 2024 (Arguello et al., 2024), and in this work, we have released 450 synthetic queries (150 per domain) from the full set of generated queries. Each query in the dataset is accompanied by its corresponding Wikidata ID, domain name, and entity name, ensuring clear entity association for retrieval experiments and analysis. Alongside these queries, we provide the source code for query generation and experimentation, as well as the visual stimuli entity set with corresponding image URLs and Wikidata ID. We also release the MTurk-based human query collection interface, allowing researchers to replicate or extend the human TOT query elicitation process.
Availability. At the time of review, LLM-elicited queries are publicly available as part of the TREC 2024 TOT track (Arguello et al., 2024) test collection and can also be accessed at the track website. [^11] https://github.com/kimdanny/llm-tot-query-elicitation The human-elicited queries will be released as part of the TREC 2025 TOT track, aligning with the SIGIR 2025 conference. Although the human-elicited queries are not yet publicly available, we have released the source code for the human query collection interface used in pilot testing on Amazon MTurk. [^12] https://github.com/kimdanny/human-tot-query-elicitation-mturk This allows researchers to explore and reproduce the query elicitation process for future development.
@@
Limitation and Future Work. A limitation of our current LLM-elicitation method is that prompts are domain-specific, limiting their generalizability. Future work should develop generalized prompting strategies to elicit TOT queries across diverse search contexts without extensive manual tuning. Additionally, expanding the methodology to multi-turn or multi-modal interactions (CH-Wang et al., 2025) could better reflect real-world TOT search behavior, where users iteratively refine queries as they recover missing information.
@@ 7. Conclusion
Acknowledgements.
We thank Ian Soboroff for implementing the in-house human query collection interface based on the MTurk pilot, and for coordinating contractor hiring and oversight at NIST. We also appreciate Alfredo Gomez for helping collect validation queries for the Landmark and Person domains, and thank Alfredo Gomez, Athiya Deviyani, Jessica Huynh, and Shaily Bhatt for their feedback during pilot testing.
@@ Appendix C Prompt for Linguistic Similarity Test in the Movie Domain
Given a paragraph, you will annotate the paragraph with the following coding scheme. Read the paragraph and annotate each sentence following the coding scheme. When outputting the response, follow the output JSON format. Coding scheme: movie: Sentence that describes characteristics of the movie itself (e.g., category, genre, tone, plot). context: Context where the searcher previously engaged with the movie. References to when, where the searcher watched the movie, who they watched it with, and even world events that were happening around the time. previous-search: Describes a previous attempt to find the movie title. References to failed search attempts, descriptions of search strategies that were unsuccessful. social: Communicates a social nicety to the community of people (e.g., If you could help I’d really appreciate it). uncertainty: Conveys uncertainty about information described (e.g., It was a foreign film I think either French or German, but I could be wrong). opinion: Conveys an opinion or judgment about some aspect of the movie (e.g., Its pretty confusing all the way to the end when there’s only one surviving woman). emotion: Conveys how the movie made the viewer feel (e.g., It was the first movie that kept me awake at night). relative-comparison: Statement that draws one or more comparisons that require additional information in order to extract the meaning of the statement (e.g., the main character looks like Brad Pitt). Paragraph: {Paragraph} Output JSON format: { ”¡Sentence Number¿”: [”¡code¿”], ”¡Sentence Number¿”: [”¡code¿”, ”¡code¿”, …], ”¡Sentence Number¿”: [] } For each sentence number in a paragraph, you can put multiple appropriate code. However, if none of the codes match with a sentence, you can leave the list empty.